OpenStudy (anonymous):

Simplify secx times cosx div1+sinx

OpenStudy (anonymous):

can anyone please help to solve this

myininaya (myininaya):

\[secx*\frac{cosx}{1+sinx}=\frac{1}{cosx}*\frac{cosx}{1+sinx}=\frac{1}{1+sinx}*\frac{1-sinx}{1-sinx}=\frac{1-sinx}{1-\sin^2x}\] =\[\frac{1-sinx}{\cos^2x}=\frac{1}{\cos^2x}-\frac{sinx}{\cos^2x}=\sec^2x-tanxsecx=secx(1-tanx)\]

myininaya (myininaya):

r we suppose to show it equals something?

OpenStudy (anonymous):

no we r suppose to show its simple form and the answer might be tanx

myininaya (myininaya):

hmmm i got secx(1-tanx) not tanx

myininaya (myininaya):

who said it might be tanx? that is wrong my simplification above looks flawless

OpenStudy (anonymous):

yeah your answer is flawless it is possible that the question is wrong

myininaya (myininaya):

did i write the question down correctly?

OpenStudy (anonymous):

\[\sec x \times (\cos x/1+\sin x)\]

myininaya (myininaya):

1+sinx is in the denominator not just 1 right?

OpenStudy (anonymous):

right